# Stability of optimal transport on metric measure spaces

### v1

@@ 2.4 Global concavity estimate
Since $a_{1}\mathfrak{m}\leq\rho\leq a_{2}\mathfrak{m}$, by [43, Corollary 2.4], $\rho$ is a doubling measure. Moreover, since $S$ is a John domain, by [26, Proposition 3.7], $\rho$ satisfies the Boman chain condition (see Definition A.2) and we can choose a covering $\mathcal{F}$, such that for any $B\in\mathcal{F}$, $r_{B}\leq 1$. Then
@@ 2.5 Stability estimate for Kantorovich functionals
$\leq$ $C_{6}\left|\mathbb{E}_{\mu^{t}[\varphi_{*}]-\mu^{t}[\psi_{*}]}(\varphi_{*}-\psi_{*})\right|^{\frac{1}{2}},$
where $C_{6}$ depends on $K,N,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S,\mathop{\rm Lip}\nolimits(\varphi),\mathop{\rm Lip}\nolimits(\psi)$.
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)=$ $-\left(\mathbb{E}_{\mu_{x}^{t}}[\phi_{s}]-\mathbb{E}_{\mu^{t}}[\phi_{s}]\right).$ (2.35)
@@
$\int_{S}\left|\Big(\Phi_{t}[\varphi_{*}](x)-\mathrm{K}_{t}[\varphi_{*}]\Big)-\Big(\Phi_{t}(\psi_{*})(x)-\mathrm{K}_{t}[\psi_{*}]\Big)\right|\,{\mathrm{d}}\rho(x)$
$\leq$ $\int_{0}^{1}\int_{S}\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s=\int_{0}^{1}\int_{S}\left|\mathbb{E}_{\mu_{x}^{t}}[\phi_{s}]-\mathbb{E}_{\mu^{t}}[\phi_{s}]\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s$
@@ 2.6 Passing to the limit
For $\delta>0$ and $Y_{\delta}:=\{y\in X:{\mathrm{d}}(y,Y)\leq\delta\}$, we have
@@
Combining with the stochastic completeness of the heat flow and the heat kernel estimate (2.10), we can find $\delta>0$ such that
@@
Since $Y$ is compact, there exists $y_{1}\in{Y_{\delta}\setminus Y}$ and $y_{2}\in Y$, such that
@@
Step 2: By uniqueness of optimal transport $T$ and (2.40), for almost every $x\in S$, there is a unique $T(x)\in Y_{\delta}$ so that
@@
Thus by the same argument as (2.44) and (2.46), for any $\epsilon>0$ we can find $0<r_{1}<r_{2}<\delta\wedge\epsilon$ such that
@@
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}\leq{e^{\frac{-\varphi^{c}(x)-3\epsilon}{t}}},~~~\forall y\in Y_{\delta}\setminus B\big(T(x),r_{2}\big),$ (2.52)
and for $t$ small enough:
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}>{e^{\frac{-\varphi^{c}(x)-2\epsilon}{t}}},~~~\forall y\in B\big(T(x),r_{1}\big).$ (2.53)
Thus
@@
Let $\varphi$ and $\psi$ be Kantorovich potentials from $\mu$ to $\rho$ and $\nu$ to $\rho$ respectively. By the heat kernel lower bound (2.10) and the stochastic completeness of the heat flow, we can see that $|\Phi_{t}(\psi_{*})|,|\Phi_{t}(\varphi_{*})|$ are uniformly bounded for $t\in(0,1]$. So by the dominated convergence theorem and Lemma 2.12 we obtain
@@
$\int_{S}\left|\Big(\varphi^{c}(x)-\mathbb{E}_{\rho}(\varphi^{c})\Big)-\Big(\psi^{c}(x)-\mathbb{E}_{\rho}(\psi^{c})\Big)\right|\,{\mathrm{d}}\rho(x)\leq\sqrt{C\mathbb{E}_{\mu-\nu}(\varphi-\psi)}$ (2.57)
@@ 3 Stability of optimal transport maps
Let $(X,{\mathrm{d}})$ be an Alexandrov space with no boundary. Then under the same assumptions of Theorem 1.4, there exists a constant $\bar{C}>0$, depending on $k,n,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),{\rm{Per}}(S),S$, such that for any $\mu,\nu\in\mathcal{P}_{2}(Y)$, we have
@@
The Riemannian structure on the set of regular points $X_{\mathrm{reg}}$ (cf.[38]) endows $\mathrm{T}X_{\mathrm{reg}}$ with the structure of a Euclidean vector bundle over $X_{\mathrm{reg}}$. On this Euclidean vector bundle, one can define (cf. [25, Section3] ) a canonical Liouville measure $\mathfrak{m}_{L}$: it is the unique Borel measure on $\mathrm{T}X_{\mathrm{reg}}$ such that for any Borel set $A\subset\mathrm{T}X_{\mathrm{reg}}$, we have
$\mathfrak{m}_{L}(A)=\int_{X}\mathcal{H}^{n}\big(A\cap\mathrm{T}_{x}X\big)\,{\mathrm{d}}\mathcal{H}^{n}(x).$
We extend $\mathfrak{m}_{L}$ to a measure on $\mathrm{T}X$ by setting $\mathfrak{m}_{L}(\mathrm{T}X\setminus\mathrm{T}X_{\mathrm{reg}})=0$. Moreover, by [6, Theorem 1.4], if $(X,{\mathrm{d}})$ is an Alexandrov space with no boundary, then the geodesic flow preserves the Liouville measure.
Let $\mathrm{S}X=\{v\in\mathrm{T}X:|v|=1\}$ denote the unit tangent bundle. There is a canonical Liouville measure $\mathfrak{m}_{S}$ on $\mathrm{S}X$ (cf.[25, Section 3.6]), also called Liouville measure, such that if the geodesic flow defined on ${\rm T}X$ preserves $\mathfrak{m}_{L}$, then the geodesic flow defined on ${\rm S}X$ preserves $\mathfrak{m}_{S}$ as well. For any Borel set $A\subseteq{\rm S}X$, we have
$\mathfrak{m}_{S}(A)=\int_{{\rm S}X}\chi_{A}(x,v)\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)=c_{n}\int_{X}\int_{\Sigma_{x}(X)}\chi_{A}(x,v)\,{\mathrm{d}}\sigma_{x}(v)\,{\mathrm{d}}\mathcal{H}^{n}(x),$ (3.2)
where $\Sigma_{x}(X)$ denotes the space of directions at $x$, $c_{n}=\mathcal{H}^{n-1}({\rm S}^{n-1})$ and $\sigma_{x}\in\mathcal{P}(\Sigma_{x}(X))$ is the canonical probability measure on the fiber.
@@
Let $\varphi_{s}:{\rm S}X\rightarrow{\rm S}X$ be the geodesic flow at time $s\in[0,1]$ on $\mathrm{S}X$, and write $\varphi_{s}(x,v)=(b_{s}(x,v),t_{s}(x,v))$, where $b_{s}(x,v)\in X$ and $t_{s}(x,v)\in{\Sigma_{b_{s}(x,v)}(X)}$. For $\mathfrak{m}_{S}$-a.e. $(x,v)\in{\rm S}X$ such that the curve $0,1]\ni s\mapsto b_{s}(x,v)$ is a locally shortest path, denote by $I_{S}(x,v)$ the set of connected components of $\{s\in[0,1]:b_{s}(x,v)\in S\}$. Since $S$ is an open set, we have
@@
For $x\in S\cap X_{\mathrm{reg}}$, by [8, Theorem 10.8.4], $\Sigma_{x}(X)$ is isometric to $S^{n-1}$, then we have
@@
$=$ $\int_{{\rm S}^{n-1}}\langle\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x),v\rangle^{2}\,{\mathrm{d}}\sigma_{x}(v)$
$=$ $d_{n}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2},$
for some universal constant $d_{n}$.
By [25, Section 3.6], $\mathfrak{m}_{S}$ is preserved by $\varphi_{s}$, so
@@
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\rm S}X}\int_{0}^{1}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\chi_{S}(b_{s}(x,v))\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)$
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\rm S}X}\sum_{i\in I_{S}(x,v)}\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{S}(x,v).$
@@
Denote $S_{1}=\{x\in X:{\mathrm{d}}(x,S)\leq 1\}$. We note that in the last integral of (3.4), only the elements $(x,v)\in{\rm S}X$ for which $x\in S_{1}$ have a non-vanishing contribution. For $\mathfrak{m}_{S}$-a.e. $(x,v)\in{\rm{S}}X$ with $x\in S_{1}$, the curve $s\mapsto b_{s}(x,v)$ is locally minimizing on $[0,1]$. Thus, applying Ohta’s semiconcavity estimate [37, Lemma 3.2] on sufficiently small subsegments yields local concavity of $u^{(x,v)}_{\mu}-\zeta|s|^{2}$ with a uniform constant $\zeta$. The local-to-global principle for concave functions then leads to the following lemma.
@@
There exists $\zeta$, depending on $k$ and $\mathop{\rm diam}\nolimits(S\cup Y)$, such that for $\mathfrak{m}_{S}$-a.e. $(x,v)\in\rm{S(S_{1})}$, the functions $u_{\mu}^{(x,v)}-\zeta|s|^{2}$, $u_{\nu}^{(x,v)}-\zeta|s|^{2}$ are concave on $s\in[0,1]$. Moreover, the modulus of their derivatives (which exist a.e. on $[0,1]$) is bounded above by $\mathop{\rm diam}\nolimits(S\cup Y)+2\zeta$.
@@
Therefore, applying Lemma 3.3 to the functions $\zeta|s|^{2}-u_{\mu}^{(x,v)}$ and $\zeta|s|^{2}-u_{\nu}^{(x,v)}$ on each compact segment $\alpha_{i}(x,v),\beta_{i}(x,v)]$ and using Hölder inequality, we obtain
[…]
$\leq$ $C_{1}\int_{{\rm S}X}\sum_{i\in I_{S}(x,v)}\left(\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}|\phi_{\mu}(b_{s}(x,v))-\phi_{\nu}(b_{s}(x,v))|^{2}\,{\mathrm{d}}s\right)^{\frac{1}{3}}\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)$
$\leq$ $C_{1}\int_{{\rm S}X}\left(\#I_{S}\right)^{\frac{2}{3}}\left(\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\right)^{\frac{1}{3}}{\mathrm{d}}\mathfrak{m}_{S}$
$\leq$ $C_{1}\left(\int_{{\rm S}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{S}\right)^{\frac{2}{3}}\left(\int_{{\rm S}X}\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{S}\right)^{\frac{1}{3}}$
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\rm S}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{S}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}},$
where $C_{1}$ depends on $k,n,a_{2},\mathop{\rm diam}\nolimits(S\cup Y)$, and the last inequality follows from the invariance of $\mathfrak{m}_{S}$ under the geodesic flow.
@@
We now aim to control $\int_{{\rm S}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{S}$, the average number of times a geodesic crosses the boundary of $S$. The argument of [26, Proposition 4.3] relies on the existence of $T>0$, so that for $(x,v)\in{\rm S}X$, the geodesic $s\mapsto b_{s}(x,v)$ is minimizing on $[0,T]$, and hence in particular does not self-intersect. In an Alexandrov space, one cannot in general expect such a uniform injectivity radius bound. We use a localization method to overcome this difficulty.
@@
$\int_{{\rm S}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)\leq c_{n}\big(\mathcal{H}^{n}(S_{1})+{\rm{Per}}(S)\big).$
@@
Let $(x,v)\in{\rm S}X$ be such that $s\mapsto b_{s}(x,v)$ is well-defined on $[0,1]$, and let $u\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$. Note that $s\mapsto b_{s}(x,v)$ has unit speed, then
@@
$\int_{{\rm S}X}\int_{0}^{1}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)$ (3.8)
$=$ $\int_{0}^{1}\int_{{\rm S}X}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)\,{\mathrm{d}}s$
@@
So up to taking a subsequence, it holds that for $\mathfrak{m}_{S}$-a.e. $(x,v)\in{\rm S}X$,
@@
$\int_{{\rm S}X}|{\mathrm{D}}\chi_{S}(b_{s}(x,v))|([0,1])\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)$ (3.10)
@@
where ()(*) follows from (3.6) and the invariance of $\mathfrak{m}_{S}$ under the geodesic flow.
@@
and that only the elements $(x,v)\in{\rm S}X$ for which $x\in S_{1}$ have a non-vanishing contribution. Combining with (3.10), we obtain
$\int_{{\rm S}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{S}(x,v)\leq$ $c_{n}\mathcal{H}^{n}(S_{1})+\frac{1}{2}c_{n}{\rm{Per}}(S).$ (3.12)
@@
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\rm S}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{S}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}}$
@@
${\mathrm{d}}\big(T_{\mu}(x),T_{\nu}(x)\big)\leq C_{2}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|,$ (3.14)
for some constant $C_{2}>0$ which depends on $k,\mathop{\rm diam}\nolimits(S\cup Y)$.

### v2

@@ 2.4 Global concavity estimate
[…]
@@ 2.5 Stability estimate for Kantorovich functionals
$\leq$ $C_{5}\left|\mathbb{E}_{\mu^{t}[\varphi_{*}]-\mu^{t}[\psi_{*}]}(\varphi_{*}-\psi_{*})\right|^{\frac{1}{2}},$
where $C_{5}$ depends on $K,N,a_{1},a_{2},\mathop{\rm diam}\nolimits(S\cup Y),S,\mathop{\rm Lip}\nolimits(\varphi),\mathop{\rm Lip}\nolimits(\psi)$.
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)=$ $-\left(\mathbb{E}_{\mu_{x}^{t}[\phi_{s}]}(v)-\mathbb{E}_{\mu^{t}[\phi_{s}]}(v)\right).$ (2.35)
@@
$\int_{S}\left|\Big(\Phi_{t}[\varphi_{*}](x)-\mathrm{K}_{t}[\varphi_{*}]\Big)-\Big(\Phi_{t}[\psi_{*}](x)-\mathrm{K}_{t}[\psi_{*}]\Big)\right|\,{\mathrm{d}}\rho(x)$
$\leq$ $\int_{0}^{1}\int_{S}\left|\frac{{\mathrm{d}}}{{\mathrm{d}}s}\Big({\Phi}_{t}[\phi_{s}]-\mathrm{K}_{t}[\phi_{s}]\Big)\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s=\int_{0}^{1}\int_{S}\left|\mathbb{E}_{\mu_{x}^{t}[\phi_{s}]}(v)-\mathbb{E}_{\mu^{t}[\phi_{s}]}(v)\right|\,{\mathrm{d}}\rho(x)\,{\mathrm{d}}s$
@@ 2.6 Passing to the limit
For $\delta>0$ and $Y_{\delta}:=\{y\in X:{\mathrm{d}}(y,Y)\leq\delta\}$, we denote
@@
Combining with the stochastic completeness of the heat flow and the heat kernel estimate (2.10), we can find $\delta>0$ large enough such that
@@
Since $Y$ is compact, there exists $y_{1}\in\overline{Y_{\delta}\setminus Y}$ and $y_{2}\in Y$, such that
@@
Step 2: By uniqueness of optimal transport $T$ and (2.40), for almost every $x\in S$, there is a unique $T(x)\in Y$ so that
@@
Then for any $\epsilon>0$, we can find $0<r_{2}<\delta$ such that
@@
$\varphi_{*}(y)-c(x,y)<-\varphi^{c}-4\delta_{r_{2}},~~~~~\forall y\in Y_{\delta}\setminus B\big(T(x),r_{2}\big)$ (2.52)
for some $\delta_{r_{2}}>0$. By Lemma 2.1, there is $t_{0}>0$ such that
$|c_{t}(x,y)-c(x,y)|<\delta_{r_{2}},~~~\forall(x,y)\in S\times Y~\text{and}~t<t_{0}.$
By the same argument as (2.44), for $t<t_{0}$ we have
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}\leq{e^{\frac{-\varphi^{c}(x)-3\delta_{r_{2}}}{t}}},~~~\forall y\in Y_{\delta}\setminus B\big(T(x),r_{2}\big),$ (2.53)
Furthermore, similar to (2.46), we can find $r_{1}<r_{2}$ so that
${e^{\frac{\varphi_{*}(y)}{t}}p_{\frac{t}{2}}(x,y)}>{e^{\frac{-\varphi^{c}(x)-2\delta_{r_{2}}}{t}}},~~~\forall y\in B\big(T(x),r_{1}\big)~\text{and}~t<t_{0}.$ (2.54)
It follows from (2.53), (2.54) that
@@
Let $\varphi$ and $\psi$ be Kantorovich potentials from $\mu$ to $\rho$ and $\nu$ to $\rho$ respectively. By the heat kernel lower bound (2.10) and the stochastic completeness of the heat flow, we can see that $|\Phi_{t}[\psi_{*}]|,|\Phi_{t}[\varphi_{*}]|$ are uniformly bounded for $t\in(0,1]$. So by the dominated convergence theorem and Lemma 2.12 we obtain
@@
$\int_{S}\left|\Big(\varphi^{c}(x)-\mathbb{E}_{\rho}(\varphi^{c})\Big)-\Big(\psi^{c}(x)-\mathbb{E}_{\rho}(\psi^{c})\Big)\right|\,{\mathrm{d}}\rho(x)\leq\sqrt{C\left|\mathbb{E}_{\mu-\nu}(\varphi-\psi)\right|}$ (2.58)
@@ 3 Stability of optimal transport maps
[…]
@@
The Riemannian structure on the set of regular points $X_{\mathrm{reg}}$ (cf.[38]) endows the tangent bundle $\mathrm{T}X_{\mathrm{reg}}$ an Euclidean vector bundle structure. On this Euclidean vector bundle, one can define (cf. [25, Section3] ) a canonical Liouville measure $\mathfrak{m}_{\mathrm{L}}$: it is the unique Borel measure on $\mathrm{T}X_{\mathrm{reg}}$ such that for any Borel set $A\subset\mathrm{T}X_{\mathrm{reg}}$, we have
$\mathfrak{m}_{\mathrm{L}}(A)=\int_{X}\mathcal{H}^{n}\big(A\cap\mathrm{T}_{x}X\big)\,{\mathrm{d}}\mathcal{H}^{n}(x).$
We extend $\mathfrak{m}_{\mathrm{L}}$ to a measure on the tangent bundle $\mathrm{T}X$ by setting $\mathfrak{m}_{\mathrm{L}}(\mathrm{T}X\setminus\mathrm{T}X_{\mathrm{reg}})=0$. Moreover, by [6, Theorem 1.4], if $(X,{\mathrm{d}})$ is an Alexandrov space with no boundary, then the geodesic flow preserves the Liouville measure.
Let $\mathrm{S}X=\{v\in\mathrm{T}X:|v|=1\}$ denote the unit tangent bundle (sphere bundle). There is a canonical Liouville measure $\mathfrak{m}_{\mathrm{S}}$ on $\mathrm{S}X$ (cf.[25, Section 3.6]), also called Liouville measure, such that if the geodesic flow defined on ${\rm T}X$ preserves $\mathfrak{m}_{\mathrm{L}}$, then the geodesic flow defined on ${\mathrm{S}}X$ preserves $\mathfrak{m}_{\mathrm{S}}$ as well. For any Borel set $A\subseteq{\mathrm{S}}X$, we have
$\mathfrak{m}_{\mathrm{S}}(A)=\int_{{\mathrm{S}}X}\chi_{A}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)=c_{n}\int_{X}\int_{\Sigma_{x}(X)}\chi_{A}(x,v)\,{\mathrm{d}}\sigma_{x}(v)\,{\mathrm{d}}\mathcal{H}^{n}(x),$ (3.2)
where $\Sigma_{x}(X)$ denotes the space of directions at $x$, $c_{n}=\mathcal{H}^{n-1}({\mathrm{S}}^{n-1})$ and $\sigma_{x}\in\mathcal{P}(\Sigma_{x}(X))$ is the canonical probability measure on the fiber.
@@
Let $\varphi_{s}:{\mathrm{S}}X\rightarrow{\mathrm{S}}X$ be the geodesic flow at time $s\in[0,1]$ on $\mathrm{S}X$, and write $\varphi_{s}(x,v)=(b_{s}(x,v),t_{s}(x,v))$, where $b_{s}(x,v)\in X$ and $t_{s}(x,v)\in{\Sigma_{b_{s}(x,v)}(X)}$. For $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$ such that the curve $0,1]\ni s\mapsto b_{s}(x,v)$ is a locally shortest path, denote by $I_{S}(x,v)$ the set of connected components of $\{s\in[0,1]:b_{s}(x,v)\in S\}$. Since $S$ is an open set, we have
@@
For $x\in S\cap X_{\mathrm{reg}}$, by [8, Theorem 10.8.4], $\Sigma_{x}(X)$ is isometric to $\mathrm{S}^{n-1}$, then there is a universal constant $d_{n}$ such that
@@
$=$ $\int_{{\mathrm{S}}^{n-1}}\langle\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x),v\rangle^{2}\,{\mathrm{d}}\sigma_{x}(v)$
$=$ $d_{n}|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|^{2}.$
By [25, Section 3.6], $\mathfrak{m}_{\mathrm{S}}$ is preserved by $\varphi_{s}$, so
@@
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\mathrm{S}}X}\int_{0}^{1}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\chi_{S}(b_{s}(x,v))\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$
$=$ $\frac{a_{2}}{c_{n}d_{n}}\int_{{\mathrm{S}}X}\sum_{i\in I_{S}(x,v)}\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}\langle\nabla\phi_{\mu}(b_{s}(x,v))-\nabla\phi_{\nu}(b_{s}(x,v)),t_{s}(x,v)\rangle^{2}\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v).$
@@
Denote $S_{1}=\{x\in X:{\mathrm{d}}(x,S)\leq 1\}$. Note that in the last integral of (3.4), only the elements $(x,v)\in{\mathrm{S}}X$ for which $x\in S_{1}$ have a non-vanishing contribution. For $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$ with $x\in S_{1}$, the curve $s\mapsto b_{s}(x,v)$ is locally minimizing on $[0,1]$. Thus, applying Ohta’s semiconcavity estimate [37, Lemma 3.2] on sufficiently small subsegments yields local concavity of $u^{(x,v)}_{\mu}-\zeta|s|^{2}$ with a uniform constant $\zeta$. The local-to-global principle for concave functions then leads to the following lemma.
@@
There exists $\zeta$, depending on $k$ and $\mathop{\rm diam}\nolimits(S\cup Y)$, such that for $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}{S_{1}}$, the functions $u_{\mu}^{(x,v)}-\zeta|s|^{2}$, $u_{\nu}^{(x,v)}-\zeta|s|^{2}$ are concave on $s\in[0,1]$. Moreover, the modulus of their derivatives (which exist a.e. on $[0,1]$) is bounded above by $\mathop{\rm diam}\nolimits(S\cup Y)+2\zeta$.
Applying Lemma 3.3 below to the functions $\zeta|s|^{2}-u_{\mu}^{(x,v)}$ and $\zeta|s|^{2}-u_{\nu}^{(x,v)}$ on each compact segment $\alpha_{i}(x,v),\beta_{i}(x,v)]$ and using Hölder inequality, we obtain
[…]
$\leq$ $C_{1}\int_{{\mathrm{S}}X}\sum_{i\in I_{S}(x,v)}\left(\int_{\alpha_{i}(x,v)}^{\beta_{i}(x,v)}|\phi_{\mu}(b_{s}(x,v))-\phi_{\nu}(b_{s}(x,v))|^{2}\,{\mathrm{d}}s\right)^{\frac{1}{3}}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$
$\leq$ $C_{1}\int_{{\mathrm{S}}X}\left(\#I_{S}\right)^{\frac{2}{3}}\left(\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\right)^{\frac{1}{3}}{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}$
$\leq$ $C_{1}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{{\mathrm{S}}X}\int_{0}^{1}|\phi_{\mu}(b_{s})-\phi_{\nu}(b_{s})|^{2}\chi_{S}(b_{s})\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{1}{3}}$
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}},$
where $C_{1}$ depends on $k,n,a_{2},\mathop{\rm diam}\nolimits(S\cup Y)$, and the last inequality follows from the invariance of $\mathfrak{m}_{\mathrm{S}}$ under the geodesic flow.
@@
We now aim to control $\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}$, the average number of times a geodesic crosses the boundary of $S$. The argument of [26, Proposition 4.3] relies on the existence of $T>0$, so that for $(x,v)\in{\mathrm{S}}X$, the geodesic $s\mapsto b_{s}(x,v)$ is minimizing on $[0,T]$, and hence in particular does not self-intersect. In an Alexandrov space, one cannot in general expect such a uniform injectivity radius bound. We use a localization method to overcome this difficulty.
@@
$\int_{{\mathrm{S}}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\leq c_{n}\big(\mathcal{H}^{n}(S_{1})+{\rm{Per}}(S)\big).$
@@
Let $(x,v)\in{\mathrm{S}}X$ be such that $s\mapsto b_{s}(x,v)$ is well-defined on $[0,1]$, and let $u\in\mathop{\rm Lip}\nolimits(X,{\mathrm{d}})$. Note that $s\mapsto b_{s}(x,v)$ has unit speed, then
@@
$\int_{{\mathrm{S}}X}\int_{0}^{1}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}s\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$ (3.8)
$=$ $\int_{0}^{1}\int_{{\mathrm{S}}X}\left|u_{k}(b_{s}(x,v))-\chi_{S}(b_{s}(x,v))\right|\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\,{\mathrm{d}}s$
@@
So up to taking a subsequence, it holds that for $\mathfrak{m}_{\mathrm{S}}$-a.e. $(x,v)\in{\mathrm{S}}X$,
@@
$\int_{{\mathrm{S}}X}|{\mathrm{D}}\chi_{S}(b_{s}(x,v))|([0,1])\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)$ (3.10)
@@
where ()(*) follows from (3.6) and the invariance of $\mathfrak{m}_{\mathrm{S}}$ under the geodesic flow.
@@
and that only the elements $(x,v)\in{\mathrm{S}}X$ for which $x\in S_{1}$ have a non-vanishing contribution. Combining with (3.10), we obtain
$\int_{{\mathrm{S}}X}\#I_{S}(x,v)\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}(x,v)\leq$ $c_{n}\mathcal{H}^{n}(S_{1})+\frac{1}{2}c_{n}{\rm{Per}}(S).$ (3.12)
@@
$\leq$ $C_{1}\left(\frac{c_{n}}{a_{1}}\right)^{\frac{1}{3}}\left(\int_{{\mathrm{S}}X}\#I_{S}\,{\mathrm{d}}\mathfrak{m}_{\mathrm{S}}\right)^{\frac{2}{3}}\left(\int_{S}|\phi_{\mu}-\phi_{\nu}|^{2}\,{\mathrm{d}}\rho\right)^{\frac{1}{3}}$
@@
${\mathrm{d}}\big(T_{\mu}(x),T_{\nu}(x)\big)\leq c|\nabla\phi_{\mu}(x)-\nabla\phi_{\nu}(x)|,$ (3.14)
for some constant $c>0$ which depends on $k,\mathop{\rm diam}\nolimits(S\cup Y)$.